Question 3 (1 point) Consider the following equation 26 log4 * = 8 What is the value of x?
Answer #1

2^{6{log_4}^x}=8

2^{6{log_4}^x}=2^3

6{log_4}^x=3

{log_4}^x=\frac{1}{2}

x=4^{\frac{1}{2}}

x=2